### v1

@@ 1 Introduction
To understand the origin of Grover’s quadratic speedup, several geometric and dynamical interpretations have been proposed. For instance, Grover’s iteration has been characterized as a geodesic in complex projective space connecting the initial and target states [miyake2001geometric], and modeled as geodesic dynamics within an information-geometric framework under the Wigner–Yanase metric [cafaro2012grover]. More recently, an interpretation based on imaginary-time evolution has been introduced in [suzuki2025grover]. Despite these valuable insights, the design of new quantum algorithms still relies largely on specialized constructions and ad hoc reasoning. A unified optimization framework that both explains Grover’s speedups and guides the development of new algorithms has yet to be established.
Manifold (Riemannian) optimization [absil2008optimization, boumal2023introduction, hu2020brief] emerges as a compelling framework for understanding Grover’s algorithms. Indeed, manifolds arise naturally in quantum systems, and manifold-based methods have already been successfully applied to tasks such as optimizing quantum circuits [luchnikov2021qgopt, wiersema2023optimizing] and quantum state and comb tomography [hsu2024quantum, li2025quantum]. In particular, the set of quantum circuits constitutes the unitary manifold $\{U\in\mathbb{C}^{N\times N}\mid U^{\dagger}U=I\}$. Since many quantum computing tasks aim to find a circuit that prepares a target state or implements a desired transformation, they can be naturally formulated as optimization problems over this unitary manifold [wiersema2023optimizing]. This suggests that manifold optimization can potentially provide a promising and systematic tool for analyzing quantum algorithms.
@@ 2.2 Geometric tools on $\mathrm{U}(N)$
As a simple example, on the sphere manifold a retraction can be implemented by normalization: $\mathrm{R}_{x}(\eta)=\frac{x+\eta}{\|x+\eta\|}$. A natural case of retraction on $\mathrm{U}(N)$ is Riemannian exponential map, defined by
@@ 6 Conclusions
This work offers a new perspective on Grover’s search algorithm from the manifold optimization. Our key contribution is the Grover-compatible retraction, a physically implementable class of retractions. We further show that the optimization dynamics remain confined to a two-dimensional invariant subspace, rendering the algorithm classically simulatable. Using standard nonconvex optimization analysis, our approach recovers the optimal $\mathcal{O}(\sqrt{N})$ quadratic speedup, providing an optimization-theoretic explanation for this quantum advantage.
From an optimization perspective, our Grover-compatible RGA method maps each mathematical iteration directly to a quantum circuit: each update simply appends gates whose generators come from the current Riemannian gradient and whose rotation angle is the step size. This enables optimization theory to move beyond the classical setting and directly inform quantum circuit design on current hardware. This viewpoint suggests several directions for future research:
  • A systematic analysis of the optimality of different retractions. As observed in our numerical experiments, the specific choice of retraction directly impacts the pullback’s Lipschitz constant, $L_{\mathrm{Rie}}$. This constant governs the theoretical step size and the resulting complexity upper bound. An optimal retraction would be one that minimizes this Lipschitz constant, thereby yielding the sharpest possible bound and the fastest guaranteed convergence.
  • The extension of this framework to other, more advanced optimization methods. While we focused on RGA with a fixed step size (which requires knowledge of $L_{\mathrm{Rie}}$), this perspective opens the door to applying more sophisticated optimization techniques. For instance, adaptive step size rules, such as the Barzilai-Borwein (BB) method [fletcher2005barzilai, iannazzo2018riemannian]. Furthermore, exploring second-order methods, like a Riemannian Newton’s algorithm, is a compelling direction.
Some of the proof strategies in this paper were interactively explored with the assistance of the large language models, and all proofs as presented here were subsequently derived, checked, and written by the authors. A detailed methodological account of this human–AI collaboration, including the interaction protocol and verification procedures, is given in companion paper [lai2025groverAI].

### v2

@@ 1 Introduction
To understand the origin of Grover’s quadratic speedup, several geometric and dynamical interpretations have been proposed. For instance, Grover’s iteration has been characterized as a geodesic in complex projective space connecting the initial and target states [miyake2001geometric], and modeled as geodesic dynamics within an information-geometric framework under the Wigner–Yanase metric [cafaro2012grover]. More recently, an interpretation based on imaginary-time evolution (ITE) has been introduced in [suzuki2025grover].
Manifold (Riemannian) optimization [absil2008optimization, boumal2023introduction, hu2020brief] emerges as a compelling framework for understanding Grover’s algorithms. Indeed, manifolds arise naturally in quantum systems, and manifold-based methods have already been successfully applied to tasks such as optimizing quantum circuits [luchnikov2021qgopt, wiersema2023optimizing] and quantum state and comb tomography [hsu2024quantum, li2025quantum]. In particular, the set of quantum circuits constitutes the unitary manifold $\{U\in\mathbb{C}^{N\times N}\mid U^{\dagger}U=I\}$. Since many quantum tasks aim to find a circuit that prepares a target state or implements a desired transformation, they can be naturally formulated as optimization problems over this unitary manifold [wiersema2023optimizing]. A series of double-bracket quantum algorithms [gluza2024double, zander2025role, suzuki2025double, gluza2024doublediag, robbiati2024double, xiaoyue2024strategies] shows that the ITE trajectories follow a Riemannian gradient flow on the unitary manifold, induced by an associated manifold optimization problem. These works suggest that manifold optimization can potentially provide a promising and systematic tool for analyzing quantum algorithms.
@@ 2.2 Geometric tools on $\mathrm{U}(N)$
As a example, on the sphere manifold a retraction can be implemented by normalization: $\mathrm{R}_{x}(\eta)=\frac{x+\eta}{\|x+\eta\|}$. A natural case of retraction on $\mathrm{U}(N)$ is Riemannian exponential map, defined by
@@ 6 Conclusions
This work presents a new viewpoint on Grover’s search algorithm through the lens of manifold optimization. Our key contribution is the Grover-compatible retraction, a physically implementable class of retractions. We further show that the optimization dynamics remain confined to a two-dimensional invariant subspace, rendering the algorithm classically simulatable. Using standard nonconvex optimization analysis, our approach recovers the optimal $\mathcal{O}(\sqrt{N})$ quadratic speedup, providing an optimization-theoretic explanation for this quantum advantage. From an optimization standpoint, our Grover-compatible RGA method maps each mathematical iteration directly to a quantum circuit: each update appends gates generated by the current Riemannian gradient, with the rotation angle given by the step size. This provides a direct pathway for optimization theory to inform quantum circuit design on contemporary hardware. We have several directions for future research:
  • A systematic study of how different retractions affect performance. Our experiments show that the choice of retraction directly influences the pullback Lipschitz constant $L_{\text{Rie }}$, which determines the allowable step size and the resulting complexity bound. An optimal retraction would minimize this constant, yielding the sharpest guarantees and the fastest provable convergence.
  • Extending this framework to more advanced optimization methods. Although we focused on fixed-step size RGA (which depends on $L_{\text{Rie }}$ ), this viewpoint naturally invites the use of more sophisticated techniques, such as adaptive step sizes (e.g., Barzilai-Borwein [fletcher2005barzilai, iannazzo2018riemannian]) and even second-order schemes like Riemannian Newton methods.
Some of the proof strategies in this paper were interactively explored with the assistance of the large language models, and all proofs as presented here were subsequently derived, checked, and written by the authors. A detailed methodological account of this human–AI collaboration, including the interaction protocol and verification procedures, is given in the companion paper [li2025groverAI].